OpenStudy (anonymous):

Find dy/dx by implicit differentiation. 3x2 + 5xy − y2 = 5

OpenStudy (zzr0ck3r):

\(3x^2+5xy-y^2=5\implies 6x+5(y+x*\frac{dy}{dx})-2y\frac{dy}{dx}=0\)

OpenStudy (zzr0ck3r):

make sense so far?

OpenStudy (zzr0ck3r):

that middle part is the product rule for \(xy\)

OpenStudy (zzr0ck3r):

and that last part is the chain rule on \(y^2\)

OpenStudy (zzr0ck3r):

so we have \(6x+5y+5x\frac{dy}{dx}-2y\frac{dy}{dx}=0\) Now solve for \(\frac{dy}{dx}\)